# Connecting Federated ADMM to Bayes

### v1

@@ 3.3 FedLap-Cov: A New ADMM Variant from VB with Flexible Covariances
Now, we can simply plug-in the form $(q_{k},\hat{t}_{k},q_{g})$ in Eq. 6 to get the update. The main differences are, first, the inclusion of $\bm{V}_{k}$ in the client’s dual term (highlighted with red) and, second, the preconditioning with $\bm{S}_{g}$ used in the server updates,
@@
FedLap-Cov: $\begin{split}\text{FedLap-Cov: }\quad\bm{w}_{k}&\leftarrow\arg\min_{\bm{w}}\;\ell_{k}(\bm{w})+\bm{v}_{k}^{\top}\bm{w}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-\mbox{$\frac{1}{2}$}\bm{w}^{\top}\bm{V}_{k}\bm{w}}+\mbox{$\frac{1}{2}$}\|\bm{w}-\bm{w}_{g}\|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{S}_{g}}}^{2}\\ \bm{S}_{k}&\leftarrow\mathbf{H}_{k}(\bm{w}_{k})-\bm{V}_{k}+\bm{S}_{g}\end{split}$ (14)
A full derivation is in App. B. Here, $\|\bm{w}\|^{2}_{\bm{S}}=\bm{w}^{\top}\bm{S}\bm{w}$ denotes the Mahalanobis distance. The two updates are performed concurrently and we have used $\mathbf{H}_{k}(\bm{w}_{k})$ by setting $\bm{w}_{\text{old}}$ to be the most recent $\bm{w}_{k}$ before updating it. These two steps along with the update for $\bm{v}_{k},\bm{V}_{k},\bm{w}_{g},\bm{S}_{g}$ in Eq. 12 constitute our new Laplace variant, which we call FedLap-Cov. In our experiments we use a diagonal matrices ($\bm{V}_{k},\mathbf{H}_{k},\bm{S}$), which already shows improved performance.
FedLap-Cov improves several aspects of ADMM and FedLap. First, the update of $\bm{v}_{k}$ uses the weighted parameters $\bm{S}_{k}\bm{w}_{k}$ and $\bm{S}_{g}\bm{w}_{g}$, weighted by their respective precisions. As a result, the uncertainty is naturally incorporated when summing $\bm{v}_{k}$ in the global updates to get $\bm{w}_{g}$. Second, the global updates use preconditioning through $\bm{S}_{g}$ which also incorporates uncertainty in the global variables. Third, $\bm{S}_{g}$ is an accumulation of all $\bm{V}_{k}$ which captures the difference between $\bm{S}_{k}$ and $\bm{S}_{g}$. Finally, the Euclidean distance is changed to Mahalanobis distance. We expect these aspects to be useful when dealing with heterogeneity across clients. The covariances should help in automatic weighing of information, which can potentially improve both performance and convergence.
@@
Therefore, if the Hessian is positive semi-definite all $\bm{V}_{k}$ will be too, which will also imply both $\bm{S}_{g}$ and $\bm{S}_{k}$ remain positive definite. We can enforce this by using a Generalised Gauss-Newton approximation (Schraudolph, 2002; Martens, 2020) or by using recent variational algorithms that ensure a positive definite covariances (Lin et al., 2020b; Shen et al., 2024). We provide the final implemented FedLap-Cov update in Eq. 21. Similarly to ADMM, PVI and FedLap, at convergence we have $\bm{w}_{g}^{*}=\bm{w}_{k}^{*}$ and $\bm{S}_{g}^{*}=\bm{S}_{k}^{*}$.
Computation and communication cost. FedLap-Cov has communication cost $2P$ from clients to server (and server to clients), as we send both $v_{k}$ and $V_{k}$ (this is for a diagonal matrix, as in all our experiments; for a full-covariance structure the theoretical cost is $23P/2+P^{2}/$ 2). Client $k$’s computation cost is $O(N_{k}P(E+1))$ as we need an additional backward pass through the dataset to calculate $V_{k}$ in our implementation, which is a small additional cost. We could alternatively use more efficient implementations, such as online implementations, to reduce this cost (Daxberger et al., 2021; Shen et al., 2024). Overall, FedLap-Cov adds new improvements to FedLap by using a more flexible posterior distribution, and we see this empirically in Sec. 4.
@@ 3.4 FedLap-Func: Improving by including function-space information
where we note that the solution to this problem (followed by calculating the Laplacian covariance at $\bm{w}_{g}$) gives us FedLap-Cov’s server update from Eq. 12, and it therefore looks very similar to Eq. 6.
Following ideas from continual learning and knowledge adaptation (Khan and Swaroop, 2021), we improve the contribution from client $k$ by sending information in function-space, instead of only sending in weight-space. We do this over a set of inputs $\mathcal{M}_{k}$ (where $\mathcal{M}_{k}$ can be different for each client, or shared). We derive a similar equation to Daxberger et al. (2023), but through a more general form which avoids a Taylor series expansion, and allows us to adapt the theory to federated learning. Each client contribution to the global server in Eq. 15 becomes (derivation in Sec. C.1),
@@
where $\ell(y_{i},\bm{w})$ is the loss at parameters $\bm{w}$ with label $y_{i}$, $\hat{y}_{i}$ are soft labels (predictions) over inputs $i\in\mathcal{M}_{k}$ using local client weights $\bm{w}_{k}$, and $\mathbf{H}^{\mathcal{M}_{k}}$ is the (approximate) Hessian of the loss over points in $\mathcal{M}_{k}$ at weights $\bm{w}_{k}$. We see that the new weight-space term is subtracted from the function-additional function-space information: this avoids double-counting the contribution from points $i\in\mathcal{M}_{k}$. Although using this in the optimisation at the global server in Eq. 15 seems to require sending additional weight-space information ($\bm{w}_{k}$ and $\mathbf{H}^{\mathcal{M}_{k}}$), the equations simplify, allowing us to still send only $\bm{V}_{k}$ and $\bm{v}_{k}$, except that compared to earlier, these now remove the contribution from points $i\in\mathcal{M}_{k}$. We additionally send $\hat{y}_{i}$ for $i\in\mathcal{M}_{k}$, which is a very small cost when the same inputs $\mathcal{M}_{k}$ are available to the clients and the server.
@@
where $\hat{y}_{i}$ is the prediction over an input $i\in\mathcal{M}_{k}$ using the previous round’s client’s model weights $\bm{w}_{k}$, $\hat{y}_{i,g}$ is the prediction using the global server weights $\bm{w}_{g}$, and we have introduced a hyperparameter $\tau$ that upweights the function-space contribution (we might want $\tau>1$ when we have few points in $\mathcal{M}_{k}$). FedLap-Func keeps the properties of FedLap at a stationary point: $\bm{w}_{g}=\bm{w}_{k}=\bm{w}_{g}^{*}$ (at a stationary point, all the function-space terms contribute zero gradient). We also see a nice property of the algorithm: FedLap-Func recovers FedLap when $\mathcal{M}_{k}=\emptyset$. FedLap-Func can therefore be seen as improving FedLap by introducing some points in function-space.
We note that our points in $\mathcal{M}_{k}$ do not have to be points from client’s data $\mathcal{D}_{k}$, and can be publicly available (unlabelled) data, like in Federated Distillation (Li and Wang, 2019; Lin et al., 2020a; Seo et al., 2020; Wu et al., 2021). Our algorithm is related to these knowledge distillation approaches in Federated Distillation, except we also send weights. We hope future work can explore these connections further.
Computation and communication cost. The communication cost of FedLap-Func per round is almost the same as FedLap: FedLap-Func has communication cost $P+C^{2}$, where $C$ is the number of classes, because we send predictions over $C$ points (one point per class in our experiments) and each prediction vector is of length $C$ ($C\ll P$). On clients, the computation cost is now $O(N_{k}PE+MPE)$, where M is the number of points in memory ($M\ll N_{k}$, so the additional cost is small). On the server, the computation cost increases compared to FedLap: the server has to optimise an objective function instead of simply adding vectors. It is now $O(MPE_{g})$, where M is the number of points in memory and $E_{g}$ is the number of optimisation steps. This is less than the computation cost at clients (because $M\ll N_{k}$) and will typically not be a problem (typically the global server is large and has more compute than clients).
@@ 4 Experiments
We run experiments on a variety of benchmarks, (i) on tabular data and image data, (ii) using logistic regression models, multi-layer perceptrons, and convolutional neural networks, (iii) at different client data heterogeneity, and (iv) with different numbers of clients. We focus on showing that (i) our VB-derived algorithm FedLap performs comparably to the best-performing ADMM algorithm FedDyn (despite having one fewer hyperparameter to tune: no local weight-decay), (ii) including full covariances in FedLap-Cov improves performance, and (iii) including function-space information in FedLap-Func improves performance. Hyperparameters and further details are in App. E.
Datasets and splits. We learn a logistic regression model on two (binary classification) datasets: UCI Credit (Quinlan, 1987) and FLamby-Heart (Janosi et al., 1988; du Terrail et al., 2022). Details on the datasets are in App. E. We use the same heterogeneous split on UCI Credit as in previous work (Ashman et al., 2022), splitting the data into 10 clients (results with a homogeneous split of UCI Credit are in Table 4 in App. F). FLamby-Heart is a naturally-heterogeneous split consisting of data from 4 hospitals (du Terrail et al., 2022). We also train a 2-hidden layer perceptron (as in previous work (Acar et al., 2021; McMahan et al., 2016)) on MNIST (LeCun et al., 1998) and Fashion MNIST (FMNIST) (Xiao et al., 2017) with $K=10$ clients. We use a random $10\%$ of the data in FMNIST to simulate having less data, and use the full MNIST dataset. We consider both a homogeneous split and a heterogeneous split. To test having more clients, we also use a heterogenous split of (full) Fashion MNIST across 100 clients. Lastly, we train a convolutional neural network on CIFAR10 (Krizhevsky and Hinton, 2009), using the same CNN from previous work (Pan et al., 2020; Zenke et al., 2017). We split the data heterogeneously into $K=10$ clients. For all heterogeneous splits, we sample Dirichlet distributions that decide how many points per class go into each client (details in Sec. E.1). Our sampling procedure usually gives 2 clients 50% of all the data, and 6 clients have 90% of the data. Within the clients, 60-95% of client data is within 4 classes out of 10.
@@
Table 1: The average accuracy (standard deviation in parentheses) over three runs reported after a few fixed number of communication rounds. We see that FedLap performs at least as well as FedDyn in all settings, while FedLap-Cov significantly improves over FedLap and FedDyn (except for MNIST homogeneous where all methods perform equally well): FedLap-Cov’s improvement over FedDyn is reported next to FedLap-Cov’s accuracy. FedLap-Func also improves performance (especially seen on CIFAR10 and heterogeneous FMNIST). For each run, we use the average accuracy over the previous 3 rounds to account for instabilities (maximum accuracy over previous 3 rounds in Table 2), and bold the top two performing algorithms (even if their standard deviations overlap with others).
Methods and hyperparameters. We compare FedLap, FedLap-Cov and FedLap-Func with three baselines: FedAvg (McMahan et al., 2016), FedProx (Li et al., 2020) and FedDyn (Acar et al., 2021), which is the best-performing federated ADMM-style method (FedDyn performs much better than FedADMM). We fix the local batch size and local learning rate (using Adam (Kingma and Ba, 2015)) to be equal for every algorithm, and sweep over number of local epochs, the $\delta$ and $\alpha$ hyperparameters, for FedDyn the additional local weight-decay hyperparameter, and for FedLap-Func the additional $\tau$ hyperparameter. For FedLap-Func, we send one randomly-selected point per class per client to the global server in the first communication round (two points for CIFAR10), and in subsequent rounds only send predictions over these same points. This breaks the strictest requirement of not sending any client data to the global server, however, this is very few points, and it might be reasonable to send a few random points having obtained prior permission. We do not report results of FedLap-Func on FLamby-Heart because of the sensitive nature of medical data.
@@
FedLap performs at least as well as FedDyn across datasets and splits, and both are better than FedAvg and FedProx. We first compare FedLap with the other baselines. We see that FedLap performs at least as well as FedDyn on all datasets and heterogeneity levels, showing that FedLap is similarly strong as FedDyn, despite having one fewer hyperparameter to tune (see also results on a homogeneous split of UCI Credit in Table 4 in App. F, where FedLap performs better than FedDyn). We note that FedDyn’s performance is very sensitive to the value of this additional weight-decay hyperparameter (we provide an example of this hyperparameter’s importance in Sec. D.1). Additionally, we find that FedLap performs well with the optimal global weight-decay $\delta$ (we show an example of FedLap’s improved performance over FedDyn due to this property in Sec. D.2).
@@
FedLap-Func improves upon FedLap. We also see that FedLap-Func improves upon FedLap, especially in the heterogeneous splits of Fashion MNIST (10 clients) and CIFAR10. FedLap-Func often reaches better accuracies at earlier rounds, as it uses function-space information while the model is still far from optimal. We note that, in our experiments, we send function-space information over very few inputs (one randomly-chosen point per class per client, or two points with CIFAR10). Future work can look at using public unlabelled datasets to send function-space information over many more inputs, like in Federated Distillation.

### v2

@@ 3.3 FedLap-Cov: A New ADMM Variant with Flexible Covariances
Now, we can simply plug-in the forms of $(q_{k},\hat{t}_{k},q_{g})$ into Eq. 6 to get the update. The main differences are, first, the inclusion of $\bm{V}_{k}$ in the client’s dual term (highlighted with red) and, second, the preconditioning with $\bm{S}_{g}$ used in the server updates,
@@
FedLap-Cov client update: $\begin{split}\text{FedLap-Cov client update: }\quad\bm{w}_{k}&\leftarrow\arg\min_{\bm{w}}\;\ell_{k}(\bm{w})+\bm{v}_{k}^{\top}\bm{w}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-\mbox{$\frac{1}{2}$}\bm{w}^{\top}\bm{V}_{k}\bm{w}}+\mbox{$\frac{1}{2}$}\|\bm{w}-\bm{w}_{g}\|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{S}_{g}}}^{2}\\ \bm{S}_{k}&\leftarrow\mathbf{H}_{k}(\bm{w}_{k})-\bm{V}_{k}+\bm{S}_{g}\end{split}$ (14)
A full derivation is in App. B. Here, $\|\bm{w}\|^{2}_{\bm{S}}=\bm{w}^{\top}\bm{S}\bm{w}$ denotes the Mahalanobis distance. The two updates are performed concurrently and we have used $\mathbf{H}_{k}(\bm{w}_{k})$ at the most recent parameters $\bm{w}_{k}$. These two steps along with the update for $\bm{v}_{k},\bm{V}_{k},\bm{w}_{g},\bm{S}_{g}$ in Eq. 12 constitute our new Laplace variant, which we call FedLap-Cov. In our experiments we use diagonal matrices ($\bm{V}_{k},\mathbf{H}_{k},\bm{S}$), which already shows improved performance.
FedLap-Cov improves several aspects of ADMM and FedLap. First, the update of $\bm{v}_{k}$ uses the parameters $\bm{S}_{k}\bm{w}_{k}$ and $\bm{S}_{g}\bm{w}_{g}$, weighted by their respective precisions. As a result, the uncertainty is naturally incorporated when summing $\bm{v}_{k}$ in the global updates to get $\bm{w}_{g}$. Second, the global updates use preconditioning through $\bm{S}_{g}$ which also incorporates uncertainty in the global variables. Third, $\bm{S}_{g}$ is an accumulation of all $\bm{V}_{k}$ which captures the difference between $\bm{S}_{k}$ and $\bm{S}_{g}$. Finally, the Euclidean distance is changed to a Mahalanobis distance. We expect these aspects to be useful when dealing with heterogeneity across clients. The covariances should help in automatic weighing of information, which can potentially improve both performance and convergence.
@@
Therefore, if the Hessian is positive semi-definite all $\bm{V}_{k}$ will be too, which will also imply both $\bm{S}_{g}$ and $\bm{S}_{k}$ remain positive definite. We can enforce this by using a Generalised Gauss-Newton approximation (Schraudolph, 2002; Martens, 2020) or by using recent variational algorithms that ensure positive definite covariances (Lin et al., 2020b; Shen et al., 2024). We provide the final implemented FedLap-Cov update in Eq. 21. Similarly to ADMM, PVI and FedLap, at convergence we have $\bm{w}_{g}^{*}=\bm{w}_{k}^{*}$ and $\bm{S}_{g}^{*}=\bm{S}_{k}^{*}$.
FedLap-Cov has communication cost $2P$ from clients to server (and server to clients), as we send both $v_{k}$ and $V_{k}$ (this is for a diagonal matrix, as in all our experiments; for a full-covariance structure the theoretical cost is $23P/2+P^{2}/$ 2). Client $k$’s computation cost is $O(N_{k}P(E+1))$ as we need an additional backward pass through the dataset to calculate $V_{k}$ in our implementation, which is a small additional cost. We could alternatively use more efficient implementations, such as online implementations, to reduce this cost (Daxberger et al., 2021; Shen et al., 2024). Overall, FedLap-Cov adds new improvements to FedLap by using a more flexible posterior distribution, and we see this empirically in Sec. 4.
@@ 3.4 FedLap-Func: Improving by including function-space information
where we note that the solution to this problem (followed by calculating the Laplace covariance at $\bm{w}_{g}$) gives us FedLap-Cov’s server update from Eq. 12, and it therefore looks very similar to Eq. 6.
Following ideas from continual learning and knowledge adaptation (Khan and Swaroop, 2021), we improve the contribution from client $k$ by also sending information in function-space, instead of only sending in weight-space. We do this over a set of inputs $\mathcal{M}_{k}$ (where $\mathcal{M}_{k}$ can be different for each client, or shared). We derive a similar equation to Daxberger et al. (2023), but through a more general form which avoids a Taylor series expansion, and allows us to adapt the theory to federated learning. Each client contribution to the global server in Eq. 15 becomes (derivation in Sec. C.1),
@@
where $\ell(y_{i},\bm{w})$ is the loss at parameters $\bm{w}$ with label $y_{i}$, $\hat{y}_{i}$ is the soft label (prediction) over an input $i\in\mathcal{M}_{k}$ using local client weights $\bm{w}_{k}$, and $\mathbf{H}^{\mathcal{M}_{k}}$ is the (approximate) Hessian of the loss over points in $\mathcal{M}_{k}$ at weights $\bm{w}_{k}$. We see that the new weight-space term is subtracted from the additional function-space information: this avoids double-counting the contribution from points $i\in\mathcal{M}_{k}$. Although using this in the optimisation at the global server in Eq. 15 seems to require sending additional weight-space information ($\bm{w}_{k}$ and $\mathbf{H}^{\mathcal{M}_{k}}$), the equations simplify, allowing us to still send only $\bm{V}_{k}$ and $\bm{v}_{k}$, except that compared to earlier, these now remove the contribution from points $i\in\mathcal{M}_{k}$. We additionally send $\hat{y}_{i}$ for $i\in\mathcal{M}_{k}$, which is a very small cost when the same inputs $\mathcal{M}_{k}$ are available to the clients and the server.
@@
where $\hat{y}_{i}$ is the prediction over an input $i\in\mathcal{M}_{k}$ using the previous round’s client model weights $\bm{w}_{k}$, $\hat{y}_{i,g}$ is the prediction using the global server weights $\bm{w}_{g}$, and we have introduced a hyperparameter $\tau$ that upweights the function-space contribution (we might want $\tau>1$ when we have few points in $\mathcal{M}_{k}$). FedLap-Func keeps the properties of FedLap at a stationary point: $\bm{w}_{g}^{*}=\bm{w}_{k}^{*}$ (at a stationary point, all the function-space terms contribute zero gradient). We also see a nice property of the algorithm: FedLap-Func recovers FedLap when $\mathcal{M}_{k}=\emptyset$. FedLap-Func can therefore be seen as improving FedLap by introducing some points in function-space.
We note that our points in $\mathcal{M}_{k}$ do not have to be points from client’s data $\mathcal{D}_{k}$, and can be publicly available (unlabelled) data, like in Federated Distillation (Li and Fedmd, 2019; Lin et al., 2020a; Seo et al., 2020; Wu et al., 2021). Our algorithm is related to these knowledge distillation approaches in Federated Distillation, except we also send weights, thereby connecting to non-distillation approaches. We hope future work can explore these connections further.
The communication cost of FedLap-Func per round is almost the same as FedLap: FedLap-Func has communication cost $P+C^{2}$, where $C$ is the number of classes ($C\ll P$), because we send predictions over $C$ points (one point per class in our experiments) and each prediction vector is of length $C$. On clients, the computation cost is now $O((N_{k}+M)PE)$, where $M$ is the number of points in memory ($M\ll N_{k}$, so the additional cost is small). On the server, the computation cost increases compared to FedLap: the server has to optimise an objective function instead of simply adding vectors. It is now $O(MPE_{g})$, where $M$ is the number of points in memory and $E_{g}$ is the number of optimisation steps. This is less than the computation cost at clients (because $M\ll N_{k}$) and will typically not be a problem (typically the global server is large and has more compute than clients).
@@ 4 Experiments
We run experiments on a variety of benchmarks, (i) on tabular data and image data, (ii) using logistic regression models, multi-layer perceptrons, and convolutional neural networks, (iii) at different client data heterogeneity levels, and (iv) with different numbers of clients. We focus on showing that (i) our VB-derived algorithm FedLap performs comparably to the best-performing ADMM algorithm FedDyn (despite having one fewer hyperparameter to tune: no local weight-decay), (ii) including full covariances in FedLap-Cov improves performance, and (iii) including function-space information in FedLap-Func improves performance. Hyperparameters and further details are in App. E.
Models and datasets. We learn a logistic regression model on two (binary classification) datasets: UCI Credit (Quinlan, 1987) and FLamby-Heart (Janosi et al., 1988; du Terrail et al., 2022). Details on the datasets are in App. E. We use the same heterogeneous split on UCI Credit as in previous work (Ashman et al., 2022), splitting the data into 10 clients (results with a homogeneous split of UCI Credit are in Table 4 in App. F). FLamby-Heart is a naturally-heterogeneous split consisting of data from 4 hospitals (du Terrail et al., 2022). We also train a 2-hidden layer perceptron (as in previous work (Acar et al., 2021; McMahan et al., 2016)) on MNIST (LeCun et al., 1998) and Fashion MNIST (FMNIST) (Xiao et al., 2017) with $K=10$ clients. We use a random $10\%$ of the data in FMNIST to simulate having less data, and use the full MNIST dataset. We consider both a homogeneous split and a heterogeneous split. To test having more clients, we also use a heterogeneous split of (full) Fashion MNIST across 100 clients. Lastly, we train a convolutional neural network on CIFAR10 (Krizhevsky and Hinton, 2009), using the same CNN from previous work (Pan et al., 2020; Zenke et al., 2017). We split the data heterogeneously into $K=10$ clients. For all heterogeneous splits, we sample Dirichlet distributions that decide how many points per class go into each client (details in Sec. E.1). Our sampling procedure usually gives 2 clients 50% of all the data, and 6 clients have 90% of the data. Within the clients, 60-95% of client data is within 4 classes out of 10.
@@
Table 1: The average accuracy (standard deviation in parentheses) over three runs reported after fixed numbers of communication rounds. We see that FedLap performs at least as well as FedDyn in all settings, while FedLap-Cov significantly improves over FedLap and FedDyn (except for homogeneous MNIST where all methods perform equally well): FedLap-Cov’s improvement over FedDyn is reported next to FedLap-Cov’s accuracy. FedLap-Func also improves performance (especially seen on CIFAR10 and heterogeneous FMNIST). For each run, we use the average accuracy over the previous 3 rounds to account for instabilities (maximum accuracy over previous 3 rounds in Table 2), and bold the top two performing algorithms (even if their standard deviations overlap with others).
Methods and hyperparameters. We compare FedLap, FedLap-Cov and FedLap-Func with three baselines: FedAvg (McMahan et al., 2016), FedProx (Li et al., 2020) and FedDyn (Acar et al., 2021), which is the best-performing federated ADMM-style method (FedDyn performs much better than FedADMM). We fix the local batch size and local learning rate (using Adam (Kingma and Ba, 2015)) to be equal for every algorithm, and sweep over number of local epochs, the $\delta$ and $\alpha$ hyperparameters, for FedDyn the additional local weight-decay hyperparameter, and for FedLap-Func the additional $\tau$ hyperparameter. For FedLap-Func, we assume one randomly-selected point per class per client is available to the global server (two points for CIFAR10). This breaks the strictest requirement of not sharing any client data with the global server, however, this is very few points, and it might be reasonable to share a few random points having obtained prior permission. We do not report results of FedLap-Func on FLamby-Heart because of the sensitive nature of medical data.
@@
FedLap performs at least as well as FedDyn across datasets and splits, and both are better than FedAvg and FedProx. We first compare FedLap with the other baselines. We see that FedLap performs at least as well as FedDyn on all datasets and heterogeneity levels, showing that FedLap is similarly strong to FedDyn, despite having one fewer hyperparameter to tune (see also results on a homogeneous split of UCI Credit in Table 4 in App. F, where FedLap performs better than FedDyn). We note that FedDyn’s performance is very sensitive to the value of this additional weight-decay hyperparameter (we provide an example of this hyperparameter’s importance in Sec. D.1). Additionally, we find that FedLap performs well with the optimal global weight-decay $\delta$ (we show an example of FedLap’s improved performance over FedDyn due to this property in Sec. D.2).
@@
FedLap-Func improves upon FedLap. We also see that FedLap-Func improves upon FedLap, especially in the heterogeneous splits of Fashion MNIST (10 clients) and CIFAR10. FedLap-Func often reaches better accuracies at earlier rounds, as it uses function-space information while the model is still far from optimal. We note that, in our experiments, we send function-space information over very few inputs (one randomly-chosen point per class per client, or two points with CIFAR10). Future work can look at using public unlabelled datasets to send function-space information over many more inputs, like in Federated Distillation. In Sec. E.5 we show that performance improves when we send function-space information over more datapoints (on FMNIST), as expected.